OpenStudy (anonymous):

find a value of k so that the following equations are consistent: kx - 3y - 5 = 0 8x + y - 17 =0 kx + 2y -10 = 0 can k have more than one value satisfying this condition?

OpenStudy (amistre64):

kx = 3y+5 kx = -2y +10

OpenStudy (amistre64):

0 = 5y -5 5 = 5y y = 1

OpenStudy (anonymous):

but what about the 2nd equation?

OpenStudy (amistre64):

kx = 8 k = 8/x when you find x :)

OpenStudy (amistre64):

8x +1 - 17 = 0 8x -16 = 0 8x = 16 x = 2

OpenStudy (amistre64):

k = 4

OpenStudy (anonymous):

can't you just add all three equations together and solve for x in terms of k? I got x = 25/(2k+8)

OpenStudy (amistre64):

add? no, you would need to use one of them as a "control" and subtract the other 2 from it if you were going that route

OpenStudy (amistre64):

x + y = 0 2x + 3y = 0 3x +4y = 0 ??? nah

OpenStudy (anonymous):

okay that's what I confused myself on. thanks.

OpenStudy (amistre64):

and since these are lines with different slopes, you cant have another value of "k" that would work.... thats just my gut feeling